# Machine Learning Assisted Parameter Tuning on Wavelet Transform Amorphous Radial Distribution Function

### v1

@@ 1 Introduction
The RDF, rooted in classical statistical mechanics, describes the number of atoms located at a given radial distance from a chosen reference atom [7, 3, 10, 11, 12, 13, 14]. If an amorphous material has a radial density distribution $\rho(r)$, then the RDF $g(r)$ can be derived using Eq. (1). Experimentally, it can also be obtained by Fourier transforming the reduced structure factor $S_{R}(q)=q(S(q)-1)$ from X-ray diffraction data, yielding the reduced RDF in Eq. (2) [15, 16, 11, 10].
@@
Additional issues arise when using Eq. (2). If the X-ray source has a low wave vector ($8\text{\AA }^{-1}$, such as Cu K-$\alpha$ and Mo K-$\alpha$, the resulting RDF has poor resolution. Consequently, the peaks that describe the atomic distribution in the amorphous material become indistinct and difficult to interpret. To address this limitation, a new mathematical approach is required. As demonstrated in [10], the Wavelet Transform (WT) offers an effective alternative. The WT-based RDF formulation used in [10] is shown in Eq. (3).
@@
To determine the parameters $a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$ [10] used experimental $S_{R}(q)$ data from the amorphous Ge–Se sample $\text{Ge}_{0.25}\text{Se}_{0.75}$ (referred to as the binary system), obtained via X-ray diffraction using a Cu K-$\alpha$ source ($\lambda\approx 1.54\ \text{\AA }$) and transformed using Eq. (3). The initial parameters were first chosen arbitrarily and then manually tuned to match the alignment of WT-derived RDF and the experimental RDF from X-ray diffraction. The best prediction result of WT-RDF was achieved with $a=0.590$, $b=4.600$, $K_{f}=0.106$, $\tilde{C}=0.212$, and $\Lambda=-4.700$. These optimized parameters were then tested on Ag-doped Ge–Se amorphous materials $\text{Ag}_{x}(\text{Ge}_{0.25}\text{Se}_{0.75})_{100-x}$ with $x=5,10,15,20,25$ (the ternary systems). The WT-RDF remained consistent, accurately identifying the first and second peak positions with errors between 2.44% and 10.16%, while also preserving the overall RDF trend [10].
However, the RDF amplitudes obtained using the WT approach still do not fully match the experimental RDFs for each material [10]. This limitation makes the WT-based RDF model insufficient for accurately determining amorphous structures. In particular, it affects the calculation of coordination numbers (CN), the number of nearest-neighbor atoms around a chosen central atom, which depends strongly on the RDF amplitude, as indicated in Eq. (5) [8].
@@
Therefore, it was still unclear if the WT-based mathematical model was defective or if the previously selected parameters were precise. In order to address this, [17] used the Born–Mayer (BM) potential to evaluate the WT-RDF model (Eq. (3)) against RDFs derived using classical density functional theory (CDFT). Because it adequately captures short-range atomic interactions, such as ionic bonds, which are typical of amorphous materials, the BM potential was chosen [18]. The WT-RDF reliably reproduces the coefficients of the BM potential, which expresses atomic repulsion and attraction, according to results from [17]. This validates the WT-RDF model from [10] and shows that inaccurate parameter selection, rather than model defects, is the main cause of disparities in RDF amplitude.
Inspired by [17], which demonstrated that just the model parameters need to be chosen more precisely, this work suggests utilizing machine learning (ML) to tune the WT-RDF model parameters. It was selected because ML uses the calculus of variations to improve parameters for predictive models [19]. Furthermore, substantial progress in component designing and production has been made possible by machine learning [20, 21]. WT-RDF++ is an improved WT-RDF model with ML-optimized parameters. Two machine learning models, the Radial Basis Function (RBF) and Long Short-Term Memory (LSTM), are used to benchmark the performance of WT-RDF++. Finally, the main research contributions are summarized into two-fold:
@@ 2 Wavelet Transform Amorphous RDF
A mathematical framework called WT-RDF uses the Wavelet Transform (WT) to calculate the RDF of amorphous materials [10]. WT was selected primarily for two reasons. In order to overcome the poor resolution of Fourier Transform (FT) when analyzing $S_{R}(q)$ data from low-energy X-rays like Cu K-$\alpha$, it first functions as a ”mathematical microscope,” offering detailed resolution in local regions of a function through its dilation ($a$) and translation ($b$) parameters [22, 23, 24]. Second, as demonstrated in Eq. (6) [25, 26], WT permits the flexible construction of wavelet functions $f(x)$ as long as they meet the usual WT criteria, which include the zero-mean and orthogonality constraints.
@@ 2.2 Application of Amorphous WT-RDF to Ge-Se (Binary) & Ag-Ge-Se (Ternaries) based amorphous solids
As mentioned in Section I, first, the existing parameters, namely $a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$, are determined arbitrarily. Then, manual tuning is performed on the binary data and the value $a=0.590$, $b=4.600$, $K_{f}=0.106$, $\tilde{C}=0.212$, and $\Lambda=-4.700$. is obtained. This value is then applied to the ternary data and the consistency of the prediction of the distances between the first and second nearest neighbors is obtained, as shown in Fig. 3.
Figure 3: Comparison between (a) experimental RDF, and (b) WT-RDF for Ge0.25Se0.75 (binary) and Agx(Ge0.25Se0.75)100-x ($x$ = 5, 10, 15, 20, 25) (ternary A to E). The parameters are $a$ = 0.590; $b$ = 4.600; $K_{f}$ = 0.106; $\Lambda$ = -4.700; $\tilde{C}$ = 0.212.
@@ 3 Method
In this research, we employ the regression method using machine learning models (i.e., Radial Basis Function (RBF) and Long Short-Term Memory (LSTM)) for benchmarking. We compare the best prediction results of the machine learning techniques with those of our proposed WT-RDF++.
@@ 3.1 Machine learning benchmark
We employ two types of machine learning models for evaluating the prediction of radial distribution data, including the Radial Basis Function (RBF) [28] and Long Short-Term Memory (LSTM) [29] models. Specifically, RBF is a feed-forward network for function approximation, while LSTM is an effective model to handle sequential data, including non-linear functions [30]. With these two distinct network characteristics, a comprehensive benchmark is established for comparative analysis against our physical model, WT-RDF++.
Table 1: Grid search on the prediction of the RBF network on binary data. The best results are highlighted in bold.
@@ 3.1.1 Radial basis function network
where $r$ is the input, $\beta$ is the width parameter to control the spread of the Gaussian function, $\hat{y}$ is the model prediction, $M$ is the number of RBF neurons, $w_{j}$ is the linear weights, $\mu_{j}$ is the center of the $j$-th, and $b$ is the bias term.
To achieve the best results, we conduct a grid search over two parameters: $\beta$ and $M$. As shown in Table 1, we can see that $\beta$ has a major influence on lowering the MAE. Based on the result, we use $M$ of 512 and $\beta$ of 1024 for comparison.
Table 2: Grid search on the prediction of the LSTM network on binary data. The best results are highlighted in bold.
@@ 3.2 WT-RDF++
To improve the predictive performance of a physics-based model with machine learning assistance, three key factors have to be considered. Those are parameter optimization, parameter bounding, and selective loss. Model learning is driven primarily by optimizing its parameters. However, solely optimizing the parameter might lead the model to get stuck in a suboptimal region of the loss surface. Therefore, a parameter bounding method is adopted to ensure that the model converges to a better solution and avoids exploding gradients where weights change drastically. Moreover, a selective loss is incorporated to guide the model toward better predictive ability on peaks.
@@ 3.2.1 Parameter optimization
$\hat{y}=\Lambda\int_{q_{min}}^{q_{max}}S(q)\biggr[\sum^{\infty}_{s=0}\beta_{s,p}\left(\dfrac{qr-b}{a}\right)^{2s+p}+\Gamma\biggr],$ (13)
@@
where $\hat{y}$ refers to $G(r)$, which is the model prediction of input data $r$. More information about Eq. (13) can be found in Section I.
@@
    $a$ controls the scaling of qrbqr-b. Smaller values of $a$ lead to a larger magnitude of the outputs, making the prediction flatter. On the contrary, higher values result in a prediction with more fluctuations.
@@
    $b$ shifts the function $qr$, making the prediction slightly move horizontally and transforming the magnitude of the wave. Additionally, the peaks are out of reach due to extreme values.
@@
    $K_{f}$ controls the amplitude of $\Phi\left(\dfrac{qr-b}{a}\right)$. This parameter affects the prediction vertically and slightly changes the width of the valley or first minimum.
@@
    $\Lambda$ scales the entire result of the integral and summation. This parameter controls the height of the peaks. Larger values result in shorter peaks, and vice versa.
@@ 3.2.3 Selective loss
where $N$ is the number of samples, $P$ is the number of peaks (two peaks in this case), $y_{n,p}$ is the ground-truth for $n$-th sample at the $p$-th selected peak bin, $\hat{y}_{n,p}$ is the model prediction at the same bin, and $m_{n,p}\in\{0,1\}$ is a binary mask (1 = this bin is a peak; 0 = not peak).
@@ 4.1 Dataset
We conduct the experiments on 6 RDF experimental dataset for amorphous $\text{Ge}_{0.25}\text{Se}_{0.75}$ as binary and $\text{Ag}_{x}(\text{Ge}_{0.25}\text{Se}_{0.75})_{100-x}$ with $x=(5,10,15,20,25)$ as ternary A to E materials. The RDF experimental data were obtained from processed data obtained using X-ray diffraction experiments sourced from Cu K-$\alpha$ with $\lambda=1.54\ \text{\AA }$ wavelength. Each dataset contains 59 pairs describing the relationship between RDF and the radial distance to an arbitrarily selected central atom.
@@ 4.2 Training implementation
All models are trained with a GPU and implemented with PyTorch. The WT-RDF++ (100 epochs) and ML models (500 epochs) are trained from scratch with the Adam optimizer [32] with a learning rate of 0.01. The ML models are pre-processed by a standard scaler or Z-score normalization [33] to stabilize the training. For WT-RDF++, we trained the model in two rounds. After finishing the first round of training, we replace the original parameters with the optimized set and modify the parameter bound values. The learning rate is decreased to 0.001 during the second stage of training to search for a better local minimum.
@@ 4.3 Evaluation metrics
To assess the prediction performance, the Mean Absolute Error (MAE), First Peak Error (FPE), and Second Peak Error (SPE) are utilized as quantitative metrics [34, 35]. A model is considered structurally accurate if peaks are also well-aligned. [35] demonstrated the sensitivity of the first and second RDF peaks to local and medium-range structure in high-density silica glass.
@@ 5.1 Ablation study
To verify the effects of parameter tuning, parameter bounding, and selective loss, an ablation analysis is performed by omitting each component. The complemented results (a)-(d) are shown in Table 3 and Fig. 4. We observe that training the parameters on WT-RDF alone can indeed lower the MAE error on average data by 38.78%. However, in Fig. 4 (a), the first and second peaks are not attained in this case. By bounding a specific sensitive parameter, the MAE, FPE, and SPE drastically improve. Since WT-RDF is a physics-based model, the radial function is naturally present. In this case, the MAE loss is replaced with a selective loss, allowing the model to focus on the first and second peaks in the data. Furthermore, employing the selective loss without parameter bounding results in a worsening of the outcome because the sensitive parameter matters, as shown in Fig. 4 (c). Specifically, when parameter $a$ is not constrained, it continues to decrease and drastically alter parameter $\tilde{C}$ in order to compensate for the impact, which worsens the results. It can also be seen that the selective loss causes the prediction to lie between the first and second peaks, meaning that the learning outcome aligns with the intended objectives. By utilizing all components, we can achieve the best result and obtain the trained parameters, improving the WT-RDF to WT-RDF++.

### v2

# Enhancing Reconstruction Capability of Wavelet Transform Amorphous Radial Distribution Function via Machine Learning Assisted Parameter Tuning

@@ 1 Introduction
The RDF, rooted in classical statistical mechanics, describes the number of atoms located at a given radial distance from a chosen reference atom [7, 3, 10, 11, 12, 13]. If an amorphous material has a radial density distribution $\rho(r)$, then the RDF $g(r)$ can be derived using Eq. (1). Experimentally, it can also be obtained by Fourier transforming the reduced structure factor $S_{R}(q)=q(S(q)-1)$ from X-ray diffraction data, yielding the reduced RDF in Eq. (2) [14, 15, 11, 10].
@@
[…]
@@
To determine the parameters $a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$ [10] used experimental $S_{R}(q)$ data from the amorphous Ge–Se sample $\text{Ge}_{0.25}\text{Se}_{0.75}$ (referred to as the binary system), obtained via X-ray diffraction using a Cu K-$\alpha$ source ($\lambda\approx 1.54\ \text{\AA }$) and transformed using Eq. (3). The initial parameters were first chosen arbitrarily and then manually tuned to match the alignment of WT-derived RDF and the Ab Initio Molecular Dynamics (AIMD) RDF simulation. The best reconstruction result of WT-RDF was achieved with $a=0.590$, $b=4.600$, $K_{f}=0.106$, $\tilde{C}=0.212$, and $\Lambda=-4.700$. These optimized parameters were then tested on Ag-doped Ge–Se amorphous materials $\text{Ag}_{x}(\text{Ge}_{0.25}\text{Se}_{0.75})_{100-x}$ with $x=5,10,15,20,25$ (the ternary systems). The WT-RDF remained consistent, accurately identifying the first and second peak positions with errors between 2.44% and 10.16%, while also preserving the overall RDF trend [10].
However, the RDF amplitudes obtained using the WT approach still do not fully match the AIMD RDF for each material [10]. This limitation makes the WT-based RDF model insufficient for accurately determining amorphous structures. In particular, it affects the calculation of coordination numbers (CN), the number of nearest-neighbor atoms around a chosen central atom, which depends strongly on the RDF amplitude, as indicated in Eq. (5) [8].
@@
Therefore, it was still unclear if the WT-based mathematical model was defective or if the previously selected parameters were precise. In order to address this, [16] used the Born-Mayer (BM) potential to evaluate the WT-RDF model (Eq. (3)) against RDFs derived using classical density functional theory (CDFT). Because it adequately captures short-range atomic interactions, such as ionic bonds, which are typical of amorphous materials, the BM potential was chosen [17]. The WT-RDF reliably reproduces the coefficients of the BM potential, which expresses atomic repulsion and attraction, according to results from [16]. This validates the WT-RDF model from [10] and shows that inaccurate parameter selection, rather than model defects, is the main cause of disparities in RDF amplitude.
To find the RDF profile of amorphous materials, researchers use simulation [18, 19] and reconstruction methods [10] using experimental data. In simulations, the methods often used are molecular dynamics [18] and density functional theory [19]. The simulation approach often involves many atoms to achieve accurate RDF profile results, while the use of reconstruction is faster because, with minimal data, the RDF profile can be known, and physical quantities related to the structure of amorphous materials, such as interatomic distances, coordination numbers, and ring statistics, can be obtained efficiently. Here, the WT-RDF method shows good reconstruction capabilities. While WT-RDF has shown promising results, it is currently less established compared to standard RDF workflows [14]. Therefore, WT-RDF, which focused on calibrated reconstruction within the Ge-Se family, is chosen as the baseline and extended in this study.
Inspired by [16], which demonstrated that just the model parameters need to be chosen more precisely, this work suggests utilizing machine learning (ML) to tune the WT-RDF model parameters. It was selected because ML uses the calculus of variations to improve parameters for reconstructive models [20]. Furthermore, substantial progress in component designing and production has been made possible by machine learning [21, 22]. WT-RDF++ is an improved WT-RDF model with ML-optimized parameters. Two machine learning models, the Radial Basis Function (RBF) and Long Short-Term Memory (LSTM), are used to benchmark the performance of WT-RDF++. Finally, the main research contributions are summarized into two-fold:
@@ 2 Wavelet Transform Amorphous RDF
A mathematical framework called WT-RDF uses the Wavelet Transform (WT) to calculate the RDF of amorphous materials [10]. WT was selected primarily for two reasons. In order to overcome the poor resolution of Fourier Transform (FT) when analyzing $S_{R}(q)$ data from low-energy X-rays like Cu K-$\alpha$, it first functions as a ”mathematical microscope,” offering detailed resolution in local regions of a function through its dilation ($a$) and translation ($b$) parameters [23]. Second, as demonstrated in Eq. (6) [24], WT permits the flexible construction of wavelet functions $f(x)$ as long as they meet the usual WT criteria, which include the zero-mean and orthogonality constraints.
@@ 2.2 Application of Amorphous WT-RDF to Ge-Se (Binary) & Ag-Ge-Se (Ternaries) based amorphous solids
As mentioned in Section I, first, the existing parameters, namely $a$, $b$, $K_{f}$, $\tilde{C}$, and $\Lambda$, are determined arbitrarily. Then, manual tuning is performed on the binary data and the value $a=0.590$, $b=4.600$, $K_{f}=0.106$, $\tilde{C}=0.212$, and $\Lambda=-4.700$. is obtained. This value is then applied to the ternary data, and the consistency of the reconstruction of the distances between the first and second nearest neighbors is obtained, as shown in Fig. 3.
Figure 3: Comparison between (a) AIMD RDF, and (b) WT-RDF for Ge0.25Se0.75 (binary) and Agx(Ge0.25Se0.75)100-x ($x$ = 5, 10, 15, 20, 25) (ternary A to E). The parameters are $a$ = 0.590; $b$ = 4.600; $K_{f}$ = 0.106; $\Lambda$ = -4.700; $\tilde{C}$ = 0.212.
Figure 4: The workflow of the proposed WT-RDF++. $r$ data and $S_{R}(q)$ versus $q$ data are used as training input to optimize the model with $a,b,\tilde{C},K_{f}$, and $\Lambda$ parameters. In the training phase, the $G(r)$ (AIMD data) is used for the loss function, minimizing the discrepancy with the new $G(r)$ (reconstruction output). The parameter bounding and selective loss were used as cost functions for training. In the inference phase, $r$ data and $S_{R}(q)$ versus $q$ data are required as the input of the trained WT-RDF++ model, resulting in the new $G(r)$.
@@ 3 Method
In this research, we employ the regression method using machine learning models (i.e., Radial Basis Function (RBF) and Long Short-Term Memory (LSTM)) for benchmarking. Compared with Fourier Transform reconstruction results for the same $q$-range of WT-RDF studied in [16], regression-based machine learning models achieve accurate RDF reconstruction. Therefore, machine learning is used as a benchmark in this study. We compare the best reconstruction results of the machine learning techniques with those of our proposed WT-RDF++. Note that the machine learning benchmark models are regressors of $r$ and $G(r)$, not scattering to RDF inversion models. Further, the workflow of the proposed WT-RDF++ is depicted in Fig. 4. Specifically, the training input data, including $r$ data and data of $q$ and $S_{R}$($q$), is fed to the model. Moreover, the parameters are optimized using the proposed parameter bounding and selective loss to improve model reconstruction. The gap between $G(r)$ of AIMD data and the reconstructed $G(r)$ is minimized using the selective loss function. In the inference phase, the $r$ data and data of $q$ and $S_{R}$($q$) are fed to the trained model, resulting in the new $G(r)$.
@@ 3.1 Machine learning benchmark
We employ two types of machine learning models for evaluating the reconstruction of radial distribution data, including the Radial Basis Function (RBF) [26] and Long Short-Term Memory (LSTM) [27] models. Specifically, RBF is a feed-forward network for function approximation, while LSTM is an effective model to handle sequential data, including non-linear functions [28]. With these two distinct network characteristics, a comprehensive benchmark is established for comparative analysis against our physical model, WT-RDF++.
Table 1: Grid search on the reconstruction of the RBF network on binary data. The best results are highlighted in bold.
@@ 3.1.1 Radial basis function network
where $r$ is the input, $\beta$ is the width parameter to control the spread of the Gaussian function, $\hat{y}$ is the model output, $M$ is the number of RBF neurons, $w_{j}$ is the linear weights, $\mu_{j}$ is the center of the $j$-th, and $b$ is the bias term.
To achieve the best results, we conduct a grid search over two parameters: $\beta$ and $M$. As shown in Table 1, we can see that $\beta$ has a major influence on lowering the Mean Absolute Error (MAE). Based on the result, we use $M$ of 512 and $\beta$ of 1024 for comparison.
Table 2: Grid search on the reconstruction of the LSTM network on binary data. The best results are highlighted in bold.
@@ 3.2 WT-RDF++
To improve the reconstructive performance of a physics-based model with machine learning assistance, three key factors have to be considered. Those are parameter optimization, parameter bounding, and selective loss. Model learning is driven primarily by optimizing its parameters. However, solely optimizing the parameter might lead the model to get stuck in a suboptimal region of the loss surface. Therefore, a parameter bounding method is adopted to ensure that the model converges to a better solution and avoids exploding gradients where weights change drastically. Moreover, a selective loss is incorporated to guide the model toward better reconstructive ability on peaks.
@@ 3.2.1 Parameter optimization
$\hat{y}=\Lambda\int_{q_{min}}^{q_{max}}S_{R}(q)\biggr[\sum^{\infty}_{s=0}\beta_{s,p}\left(\dfrac{qr-b}{a}\right)^{2s+p}+\Gamma\biggr],$ (13)
@@
where $\hat{y}$ refers to $G(r)$, which is the model reconstruction of input data $r$. More information about Eq. (13) can be found in Section I.
@@
    $a$ is a dilation coefficient controlling the scaling of qrbqr-b. Smaller values of $a$ lead to a larger magnitude of the outputs, making the reconstruction flatter. On the contrary, higher values result in a reconstruction with more fluctuations.
@@
    $b$ is a translation coefficient shifting the function $qr$, making the reconstruction slightly move horizontally and transforming the magnitude of the wave. Additionally, the peaks are out of reach due to extreme values.
@@
    $K_{f}$, coupling parameter that controls the interaction strength among the atoms. This parameter affects the reconstruction vertically and slightly changes the width of the valley or first minimum.
@@
    $\Lambda$ is used for normalization, scaling the entire result of the integral and summation. This parameter controls the height of the peaks. Larger values result in shorter peaks, and vice versa.
@@ 3.2.3 Selective loss
where $N$ is the number of samples, $P$ is the number of peaks (two peaks in this case), $y_{n,p}$ is the ground-truth for $n$-th sample at the $p$-th selected peak bin, $\hat{y}_{n,p}$ is the model output at the same bin, and $m_{n,p}\in\{0,1\}$ is a binary mask (1 = this bin is a peak; 0 = not peak).
@@ 4.1 Dataset
We use six AIMD RDF dataset taken from [10] for amorphous $\text{Ge}_{0.25}\text{Se}_{0.75}$ as binary and $\text{Ag}_{x}(\text{Ge}_{0.25}\text{Se}_{0.75})_{100-x}$ with $x=(5,10,15,20,25)$ as ternary A to E materials, respectively. Specifically, the RDF simulation data were obtained from Ab Initio Molecular Dynamics (AIMD) using the SIESTA software, where each dataset contains 58 pairs. Meanwhile, the data of the relationship between $q$ and $S_{R}$($q$) (reduced structure factor) with 148 pairs for all amorphous materials was obtained from the X-ray diffraction experiment using an X-ray source of Cu K-$\alpha$ with $\lambda=1.54\ \text{\AA }$ wavelength, also taken from [10]. The X-ray diffraction produced a $q_{\text{min}}$ value of 0.65 $\text{\AA }^{-1}$ and a $q_{\text{max}}$ value of 8.00 $\text{\AA }^{-1}$, where these values correspond to scattering angles of 4.57°and 78.8°. It can be seen that the main peak locations of $S_{R}(q)$ are consistent with the $S(q)$ simulation results from [30] in $q_{peak}=\{{1.05,2.05,3.55,5.70\}}$. Therefore, the used $S_{R}(q)$ data in this study is considered valid for input data. The data on the relationship between $q$ and $S_{R}$($q$) for all amorphous materials were corrected from background scattering, normalization, and correction steps, using the method described in [11].
@@ 4.2 Training implementation
All models are trained with a Graphics Processing Unit (GPU) and implemented with PyTorch. The WT-RDF++ (100 epochs) and Machine Learning (ML) models (500 epochs) are trained from scratch with the Adam optimizer [31] with a learning rate of 0.01. The ML models are pre-processed by a standard scaler or Z-score normalization [32] to stabilize the training. For WT-RDF++, we trained the model in two rounds. After finishing the first round of training, we replace the original parameters with the optimized set and modify the parameter bound values. The learning rate is decreased to 0.001 during the second stage of training to search for a better local minimum.
@@ 4.3 Evaluation metrics
To assess the reconstruction performance, the Mean Absolute Error (MAE), First Peak Error (FPE), and Second Peak Error (SPE) are utilized as quantitative metrics [33, 34]. A model is considered structurally accurate if peaks are also well-aligned. [34] demonstrated the sensitivity of the first and second RDF peaks to local and medium-range structure in high-density silica glass.
@@ 5.1 Ablation study
To verify the effects of parameter tuning, parameter bounding, and selective loss, an ablation analysis is performed by omitting each component. The complemented results (a)-(d) are shown in Table 3 and Fig. 5. We observe that training the parameters on Wavelet Transform Radial Distribution Function (WT-RDF) alone can indeed lower the Mean Absolute Error (MAE) on average data by 38.78%. However, in Fig. 5 (a), the first and second peaks are not attained in this case. By bounding a specific sensitive parameter, the MAE, First Peak Error (FPE), and Second Peak Error (SPE) drastically improve. Since WT-RDF is a physics-based model, the radial function is naturally present. In this case, the MAE loss is replaced with a selective loss, allowing the model to focus on the first and second peaks in the data. Furthermore, employing the selective loss without parameter bounding results in a worsening of the outcome because the sensitive parameter matters, as shown in Fig. 5 (c). Specifically, when parameter $a$ is not constrained, it continues to decrease and drastically alter parameter $\tilde{C}$ in order to compensate for the impact, which worsens the results. It can also be seen that the selective loss causes the reconstruction to lie between the first and second peaks, meaning that the learning outcome aligns with the intended objectives. By utilizing all components, we can achieve the best result and obtain the trained parameters, improving the WT-RDF to WT-RDF++.